kinetic energy of 16g of o2 gas at 300k is same as that of 4g of ch4 gas at certain temperature T 

Grade:11

1 Answers

Rituraj Tiwari
3 years ago
Kinetic energy of 16g of O₂ gas at 300K is same as that 4g of CH₄ gas at temperature T.

We have to find value of T

★ Kinetic energy of n moles of gas is given by

K = 3/2 nRT

So first of all we need to find moles of both gases.

1.Moles of oxygen gas:

⭆ n₁ = 16/32

⭆n₁=0.5

2.Moles of methane gas:

⭆ n₂ = 4/16

⭆n₂ = 0.25

ATQ,

➠ K₁ = K₂

➠ 3/2 n₁RT₁ = 3/2 n₂RT₂

➠ n₁T₁ = n₂T₂

➠ (0.50)(300) = 0.25T₂

➠T₂ = 600K